OpenStudy (anonymous):

integral of (x^2-a^2)^(3/2) dx ... ?

myininaya (myininaya):

\[\int\limits_{}^{}(x^2-a^2)^\frac{3}{2} dx\] TRig substition! :)

OpenStudy (anonymous):

a nice trig sub for myininaya, queen of the trig substitutions

OpenStudy (anonymous):

yes but how to use it ? i don't see it

OpenStudy (anonymous):

actually i think this is kind of a pain \[x=a\tan(\theta)\] or \[x=a\sec(\theta)\] probably the second one

OpenStudy (across):

It's the second one.

OpenStudy (anonymous):

x=a*cosh(u) works also

OpenStudy (anonymous):

\[x=a\sec(\theta),dx=a\sec(\theta) \tan(\theta)d\theta\] then \[\sqrt{a^2\sec^2(\theta)-a^2}=a\tan(\theta)\] so you get \[\int a^3\tan^3(\theta) \sec(\theta)\tan(\theta)d\theta\]

myininaya (myininaya):

 so we will use a sec sub. you coud use csc if you like though (rearrage your triangle) \[otherleg=\sqrt{x^2-a^2}\] \[\sec(\theta)=\frac{x}{a}\] \[\tan(\theta) \sec(theta) d\theta=\frac{1}{a} dx\] \[\int\limits_{}^{}((a sec(\theta))^2-a^2)^\frac{3}{2} a \sec(\theta) \tan(\theta) d \theta\] \[\int\limits_{}^{} (a^2)^\frac{3}{2}(\tan^2(\theta))^\frac{3}{2} \sec(\theta) \tan(\theta) d \theta\]

OpenStudy (anonymous):

i knew you had it!

myininaya (myininaya):

\[\int\limits_{}^{} a^3 \tan^3(\theta) \sec(\theta) \tan(\theta) d \theta \]

OpenStudy (anonymous):

oh you got what i got. now you still have to integrate \[a^3\int\tan^4(\theta)\sec(\theta)d\theta\]

myininaya (myininaya):

\[\int\limits_{}^{} a^3 \tan^4(\theta) \sec(\theta) d \theta \] \[ a^3 \int\limits_{}^{} (\sec^2(\theta)-1)^2 \sec(\theta) d \theta \] \[a ^3 \int\limits_{}^{}(\sec^4(\theta)-2 \sec^2(\theta) +1) \sec(\theta) d \theta\] \[a^3 \int\limits_{}^{}( \sec^5(\theta) -2 \sec^3(\theta)+\sec(\theta) ) d \theta\]

myininaya (myininaya):

so looks at \[\int\limits_{}^{}\sec(\theta) \cdot \frac{\sec(\theta)+\tan(\theta)}{\sec(\theta)+\tan(\theta)} d \theta=\int\limits_{}^{} \frac{du}{u}=\ln|u|+c_1\] \[=\ln|\sec(\theta)+\tan(\theta)|+c_1\] now we will look at \[\int\limits_{}^{}\sec^3(\theta) d \theta=\int\limits_{}^{}\sec^2(\theta) \sec(\theta) d \theta\] \[\int\limits_{}^{}\sec^3(\theta) d \theta=\tan(\theta) \sec(\theta)-\int\limits_{}^{}\tan(\theta) \tan(\theta) \sec(\theta) d \theta\] \[\int\limits_{}^{}\sec^3(\theta) d \theta=\tan(\theta) \sec(\theta)-\int\limits_{}^{}(\sec^2(\theta)-1)\sec(\theta) d \theta\] \[\int\limits_{}^{}\sec^3(\theta) d \theta=\tan(\theta) \sec(\theta) -\int\limits_{}^{}\sec^3(\theta) d \theta +\int\limits_{}^{}\sec(\theta) d \theta\] now add that integral of sec^3(theta) d theta on both sides \[2 \int\limits_{}^{}\sec^3(\theta) d \theta=\tan(\theta) \sec(\theta) +\ln|\sec(\theta)+\tan(\theta)|+c_2\]

myininaya (myininaya):

now that last part \[\int\limits_{}^{}\sec^5(\theta) d \theta=\int\limits_{}^{} \sec^3(\theta) \sec^2 (\theta)d \theta\] \[\int\limits_{}^{} \sec^3(\theta) (1+\tan^2(\theta)) d \theta\] \[\int\limits_{}^{}\sec^3(\theta) d \theta +\int\limits_{}^{}\sec^3(\theta) \tan^2(\theta) d \theta\] \[\frac{1}{2}(\tan(\theta) \sec(\theta)+\ln|\sec(\theta)+\tan(\theta)|)+\int\limits_{}^{}\sec^3(\theta) \tan^2(\theta) d \theta\]

OpenStudy (anonymous):

but you have x not theta...

OpenStudy (anonymous):

just went for a sandwich and a beer, came back and you are still hard at work...

myininaya (myininaya):

we still aren't finishing integrating yet

myininaya (myininaya):

i can only do one thing at a time

myininaya (myininaya):

you finish it i dare you

OpenStudy (anonymous):

i don't know how

myininaya (myininaya):

i need a break its long

OpenStudy (anonymous):

some people like calc two...

OpenStudy (anonymous):

this is why god invented maple

OpenStudy (anonymous):

oh and wolfram. but that gets it

myininaya (myininaya):

ok well i should have written =\[\int\limits_{}^{}\sec^5(\theta) d \theta=\int\limits_{}^{}\sec^4(\theta) \sec(\theta) d \theta\] =\[\int\limits_{}^{}(\tan^2(\theta)+1)^2 \sec(\theta) d \theta \] =\[\int\limits_{}^{}(1+2 \tan^2(\theta) +\tan^4(\theta)) \sec(\theta) d \theta \] =\[\int\limits_{}^{} \sec(\theta) d \theta +\int\limits_{}^{} 2 \tan^2( \theta) \sec(\theta) d \theta+ \int\limits_{}^{} \tan^4(\theta) \sec(\theta) d \theta\] =\[\ln|\sec(\theta)+\tan(\theta)| +2 \int\limits_{}^{} (\sec^2(\theta)-1) \sec(\theta) d \theta + \int\limits_{}^{} (\sec^2(\theta)-1)^2 \sec(\theta) d \theta \]